# Zero Trust Architecture: A Systematic Literature Review

### v1

@@ 1 Introduction
The increasing connectivity between people, systems, and organizations due to the emergence of various cutting-edge technologies such as the Internet of Things (IoT) and cloud computing has extended the traditional network boundaries leading to highly distributed IT networks with diminishing relevance of perimeter-based security models (PBSM) [1], [2]. Similarly, as the complexity of these networks increased, enterprises manage multiple internal networks while connecting to external entities including remote offices and partner networks [3]. Moreover, access to systems and network resources comes from a wider range of devices due to operational concepts such as bring your own device (BYOD) and work from home (WFH) [3]. This expanded connectivity not only increases the attack surface but also aggregate systems with varying security strengths, making system management challenging. At the same time, the threats landscape is evolving with more frequent and sophisticated cyber attacks [4]. Consequently, Zero Trust Architecture (ZTA) has been recognized by organizations as a potential solution to mitigate those issues [5], [6]. ZTA is a security strategy based on the notion of ”never trust, always verify,” where users, devices and any other entity that can be regarded as subjects to resources on the network should not be trusted implicitly until proven regardless of their network location (i.e., internal or external) [7], [8]. According to the department of defense (DoD), ZTA ”is a dramatic paradigm shift in philosophy of how we secure our infrastructure, networks, and data, from verify once at the perimeter to continual verification of each user, device, application, and transaction.” [9].
@@
ZTA is a new paradigm of access control - the ability to determine a subject’s (a verified user or a process running on that user’s behalf) privileges and restrict access appropriately [6]. It emphasizes continuous authentication and conditional authorization to mitigate evolving cyber threats [10]. Also, ZTA focuses on the principle of least privilege, where users, devices, and processes are only allowed access to certain resources on a per request or transaction basis in accordance with their access control rights and authorizations through continuously verified credentials [1], in contrast to traditional security models that grant broad access after successful authentication [5]. In other words, ZTA continuously verifies the identity of users and devices during the communication process as opposed to determining once for the whole session enabling enhanced security in highly dynamic and interconnected ecosystems [11]. This gives the system the ability to control and adjusts the security clearance required to access a particular resource [7]. Furthermore, as data-centric framework, ZTA acts under the assumption of a hostile environment, utilizing continuous authentication and dynamic authorization at every point of access to decide whether to grant access or not, and consequently, mitigate lateral movement by malicious actors and reduce insider threats [12], [13]. Through this proactive and adaptive model of security, ZTA enhances the entire security posture for an organization by mitigating the impact of compromised devices or bad actors [14].
@@
As organizations become more dependent on digital infrastructures, the multifaceted approaches by which people access corporate networks either from within or remotely using personal or corporate devices and their heterogeneity, the increasing collaboration and inter-connectivity both from individual level and organizational which can be with subsidiaries and partner networks, and the rise of advanced and evolving cyber attacks, the urgency for an effective and adaptive security framework like ZTA is apparent. According to a Gartner’s survey report, more than 50% of companies worldwide have either partially or fully implemented zero trust strategy [15]. Also, the global market value of zero trust is expected to reach 133 billion US dollars by 2032 compared to 31.63 in 2023 according to statista [16]. These figures demonstrate the growing demand for ZTA in a constantly shifting threat environment.
@@
Similarly, there is an ongoing trend of research in this area, with several studies being conducted to analyze ZTA itself and its core principles and components [5], [14], [7], [17], [18], [6], highlighting its significance [14], [7], [5], evaluating its security potentials [3], [8], ways of enhancing the ZTA framework [13], [19], [11], how to implement or transition from PBSM to the ZTA [1], [20], [21], [14], [7], [6], [22], deploying the ZTA on a given application [23], [14], [24],[12], [25], [26], [27], and proposing security architectures based on the ZTA [28], [4], [29], [30], [26], [31]. All these studies, and other publications indicate the extent to which research on ZTA has been carried out. However, there was no study that systematically review the existing literature on ZTA with an aim to explore the various domains where ZTA is applied and the associated challenges, as well as the enabling technologies that are integrated with ZTA to implement and enhance security resilience.
@@
The purpose of this research is to perform a systematic literature review (SLR) focused on the applications of ZTA. Considering the recent boom in research in this area, the conducted SLR aims to gather, and synthesise the existing studies on ZTA with a goal to provide a comprehensive taxonomy useful for researchers and practitioners who wish to apply the ZTA-oriented framework to fortify security defenses. The key contributions of this SLR are summarized as follows:
@@
  • Our study explores and synthesizes the diverse domains where ZTA has been applied, includeing concrete examples and case studies from the reviewed literatures.
  • We present a critical analysis of enabling technologies that facilitate ZTA adoption which can aid in understanding the technological landscape that supports ZTA frameworks.
@@ 2.1 ZTA Vs PBSM
Access control is a fundamental aspect of the field of cybersecurity, regarding the management of who or what can view, modify, or use resources in an information system. Access to information systems is prevented from unauthorized entities through two broad categories – Perimeter-Based Security Model (PBSM) and dynamic ZTA.
@@
Similarly, users can safely and remotely access the company network with Virtual Private Networks (VPN) [35]. VPNs make it more difficult for hackers to intercept or steal private information by encrypting communication between users’ devices and the company network. However, under these models, everything within the confines of a network is presumed secure and trustworthy, and that cyber threats mainly originate from external actors [14]. In such a setup, if the user has been granted access, then it could be considered trustworthy during its communication lifetime [7]. Yet investigations indicate that insider threats are responsible for a large percentage of cyber incidents. In addition, a single account may be compromised resulting in an escalation of privileges and presence for the attacker in the network [36] with consequences that can be severe depending on the goals of the attacker.
@@
However, the innovative approach to protect todays distributed and heterogeneous cyber landscape is Zero Trust Architecture (ZTA) [14] which is gradually gaining its foothold in the cybersecurity space. Driven by a ”never trust, always verify” philosophy, ZTA fundamentally changes how we think about network security by assuming that every part of the infrastructure is compromised and focuses on minimizing uncertainty by enforcing precise, least-privilege, per-request access decisions [5],[22]. It emphasizes situational awareness, dynamic authentication and conditional authorization, small trust zones, and continuous trust evaluation. According to a NIST special publication [37], the main goals of ZTA are data breach prevention, insider threat detection, and attacker lateral movement restriction [30]. Nowadays, the rate at which data is generated is unprecedented given the proliferation of smart devices and increasing digital transformation. This creates serious concern for both individuals and organizations when this data is disclosed or used illegally without authorization [38]. Insider threats may be malicious users, negligent users, or compromised users [21]. Because these users often have access to sensitive resources and may be well-versed in the security mechanisms in place, insider attacks pose a serious risk to organisations [39].
@@
According to a Cybersecurity Insiders report, insider attacks increased by over 47% between 2022 and 2024, affecting 34% of firms worldwide [40]. Lateral movement is a tactical exploration of a corporate network by attackers [41]. Usually, when a weak or vulnerable system (with low privileged account) is compromised, the attackers attempt to escalate their privileges while infecting other systems in the network untill a set of target systems are reached [42]. The attacker continues to have access to the systems after the lateral movement in order to eventually accomplish some malicious objective, like data exfiltration or service interruption [43]. Hence, a system of continuous monitoring that examines each access request and dynamically modifies permissions based on the perceived trustworthiness of the access attempt in order to safeguard the systems, networks, and devices against malicious traffic is required [44]. This is accomplished by continuously monitoring and analysing network traffic, user and device behaviour, and other contextual data [5]. By putting security controls at the point of access, ZT actively identifies and prevents malicious or suspected individuals from gaining access to and manipulating critical data.
@@
Among the key difference between ZTA and legacy PBSM is that, while PBSM protects the enterprise (i.e., the resources) as a whole which makes the enterprise to suffer greatly during a security breach, ZTA safeguards the individual resources [7]. It achieves this security feat by relying on some key assumptions and fundamental principles. These are contained in several articles including the study by N. Nahar et al. [14], Yuanhang et al. [18], and S. Hasan et al. [31].
@@ 2.2 ZTA Core Components
Policy Engine (PE): Within the PE, there is Zero Trust (ZT) algorithm which acts as the brain of the ZTA and makes decision based on enterprise policies to grant, deny, or restrict access to a resource on the network [6]. The ZT algorithm also receives input from external sources such as the Continuous Diagnostic and Mitigation (CDM) systems, Records of network & system activity, activity logs, threat intelligence, etc [17], [20].
Policy Administrator (PA): The PA works closely with the PE to ultimately allow or deny a session. It setup the PEP to enforce the decision or directive from the PE to allow or deny a session [6]. The PE together with the PA make up the control plane of the ZTA framework [18]
Policy Enforcement Point (PEP): The PEP dynamically interface with the entity (subject) requesting access to a resource and the resource on the network [14]. It serves as first point of contact by a request [20], and enables, monitors and terminate access between a subject and a requested resource [29].
@@ 3 History and Publication Trend
Zero Trust Architecture (ZTA) has evolved from a security concept into an essential network security solution and is increasingly recognized by governments, industry, and academia [45]. It is a security strategy that challenges the idea that users, devices, and network components should not be trusted implicitly because of where they are (i.e., their location) within a network [46]. The idea of zero trust has been around for sometime, but many articles are attributing its formal establishment to John Kindervag (2010), a research analyst working at Forrester [18], [45]. Kindervag et al. of Forrester Research [47] in a report, titled: ”No More Chewy Centers: Introducing The Zero Trust Model Of Information Security”, outlined and described four critical flaws of traditional security models that trust internal networks by default, leaving them vulnerable to insider threats and advanced attacks. In an effort to mitigate those security issues, they introduced the idea of ”Zero Trust” based on three fundamental security concepts which are: ”Ensure That All Resources Are Accessed Securely Regardless Of Location”, ”Adopt A Least Privilege Strategy And Strictly Enforce Access Control”, and ”Inspect And Log All Traffic”. The philosophy of ZTA that is currently available in the literature is based on these concepts.
@@
However, some of the earliest work on what we refer now as Zero Trust began in early 2000s with a group of CISOs known as ”The Jericho Forum” [48], [45]. The group advocated for a new security concept called de-perimeterization, which reimagined enterprise security and arose from challenges such as managing complex firewalls amidst growing traffic volume and diversity, addressing vulnerabilities within internal network domains, and adapting to evolving business practices and working arrangements [49]. Similarly, Yong et al. [50] in 2007, while highlighting the issue of security and privacy in vehicular networks, proposed an authentication protocol using random key-set. The protocol ensures user privacy based on zero-trust policy where no central entity is trusted, i.e., the vehicles do not trust public or private networks and servers. Figure  2 shows some notable events about ZTA over the years.
In the early years of ZTA, the concept received very few contributions from researchers even after the proposal of Kindervag as shown in Figure  3. The data was obtained from Scopus database using relevant keywords for researches under the domain of computer science and Engineering. These years can be characterized with foundational works on ZTA, defining its principles, primarily led by industry thought leaders before early research papers began exploring the architectural frameworks, practical implementations, and initial use cases in enterprise environments. It was in 2020 that the trend of ZTA-related research started to pick up where publications increased exponentially with each year almost doubling its preceding year. This surge in publications can be attributed to a confluence of factors that heightened the importance and urgency of robust cybersecurity measures which include:
@@
  • Adoption of Cloud Computing and BYOD: To effectively enable remote work, there was an increase in the adoption rate of cloud computing by organizations and businesses [52], [53] with employees increasingly using personal devices (BYOD) to access corporate systems [54]. This trend created diverse attack surfaces and underscores the need for enhanced security models that can effectively manage and secure distributed environments [55]. Zero Trust with its prinnciples became essential for ensuring secure cloud access and mitigating risks posed by unmanaged devices and lateral movement of attackers.
  • Government Initiatives, Policy Mandates, and Standardization: Another potential that sparked publications in ZTA is governments initiatives and policies mandating or encouraging the adoption of Zero Trust architectures to enhance national cybersecurity resilience. For example, the executive order on cybersecurity 14028 released by the U.S. federal government in May 2021 [56]. This has led to the development of specific frameworks, such as CISA’s Zero Trust Maturity Model [57]. Similarly, the release of special publication 800 - 207 on ZTA in August 2020 by National Institute of Standards and Technology (NIST) [58] may have facilitated academic research and practical implementations, leading to increased scholarly activity.
  • Escalation of Cyber Threats and High-Profile Cyber-attacks: The period between 2020 and 2023 experienced a notable increase in cyber threats which include ransomware attacks, data breaches, e.t.c. High-profile incidents highlights the inadequacies of traditional security approaches. These include the SolarWinds cyberattack in 2020, the Colonial pipeline ransomware attack in 2021, the Microsoft exchange server exploit in 2021, the MOVEit file transfer vulnerabilities 2023, e.t.c. These threats prompted organizations to adopt more resilient security architectures like Zero Trust to better protect sensitive data and critical infrastructure, and may have influence the surge in publications within this timeframe.

### v2

@@ 1 Introduction
The increasing connectivity between people, systems, and organizations due to the emergence of various cutting-edge technologies such as the Internet of Things (IoT) and cloud computing has extended the traditional network boundaries leading to highly distributed IT networks with diminishing relevance of perimeter-based security models (PBSM) [1], [2]. Similarly, as the complexity of these networks increased, enterprises manage multiple internal networks while connecting to external entities including remote offices and partner networks [3]. Moreover, access to systems and network resources comes from a wider range of devices due to operational concepts such as bring your own device (BYOD) and work from home (WFH) [3]. This expanded connectivity not only increases the attack surface but also aggregate systems with varying security strengths, making system management challenging. At the same time, the threats landscape is evolving with more frequent and sophisticated cyber attacks [4]. Consequently, Zero Trust Architecture (ZTA) has been recognized by organizations as a potential solution to mitigate those issues [5], [6]. ZTA is a security strategy based on the notion of ”never trust, always verify,” where users, devices, and any other entity that can be regarded as subjects to resources on the network should not be trusted implicitly until proven regardless of their network location (i.e., internal or external) [7], [8]. According to the Department of Defense (DoD), ZTA ”is a dramatic paradigm shift in philosophy of how we secure our infrastructure, networks, and data, from verify once at the perimeter to continual verification of each user, device, application, and transaction.” [9].
@@
ZTA is a new paradigm of access control - the ability to determine a subject’s (a verified user or a process running on that user’s behalf) privileges and restrict access appropriately [6]. It emphasizes continuous authentication and conditional authorization to mitigate evolving cyber threats [10]. Also, ZTA focuses on the principle of least privilege, where users, devices, and processes are only allowed access to certain resources on a per request or transaction basis in accordance with their access control rights and authorizations through continuously verified credentials [1], in contrast to traditional security models that grant broad access after successful authentication [5]. In other words, ZTA continuously verifies the identity of users and devices during the communication process as opposed to determining once for the whole session enabling enhanced security in highly dynamic and interconnected ecosystems [11]. This gives the system the ability to control and adjust the security clearance required to access a particular resource [7]. Furthermore, as data-centric framework, ZTA acts under the assumption of a hostile environment, utilizing continuous authentication and dynamic authorization at every point of access to decide whether to grant access or not, and consequently mitigate lateral movement by malicious actors and reduce insider threats [12], [13]. Through this proactive and adaptive security model, ZTA enhances the entire security posture for an organization by mitigating the impact of compromised devices or bad actors [14].
@@
As organizations become more dependent on digital infrastructures, the multifaceted approaches by which people access corporate networks either from within or remotely using personal or corporate devices and their heterogeneity, the increasing collaboration and interconnectivity both from individual and organizational level which can be with subsidiaries and partner networks, and the rise of advanced and evolving cyber attacks, the urgency for an effective and adaptive security framework like ZTA is apparent. According to a Gartner’s survey report, more than 50% of companies worldwide have either partially or fully implemented zero trust strategy [15]. Also, the global market value of zero trust is expected to reach 133 billion US dollars by 2032 compared to 31.63 in 2023 according to statista [16]. These figures demonstrate the growing demand for ZTA in a constantly shifting threat environment.
@@
Similarly, there is an ongoing trend of research in this area, with several studies being conducted to analyze ZTA itself and its core principles and components [5], [14], [7], [17], [18], [6], highlighting its significance [14], [7], [5], evaluating its security potentials [3], [8], ways of enhancing the ZTA framework [13], [19], [11], how to implement or transition from PBSM to the ZTA [1], [20], [21], [14], [7], [6], [22], deploying the ZTA on a given application [23], [14], [24],[12], [25], [26], [27], and proposing security architectures based on the ZTA [28], [4], [29], [30], [26], [31]. All these studies, and other publications indicate the extent to which research on ZTA has been carried out. However, there was no study that systematically review the existing literature on ZTA with the aim to explore the various domains where ZTA is applied and the associated challenges, as well as the enabling technologies that are integrated with ZTA to implement and enhance security resilience.
@@
The purpose of this research is to perform a systematic literature review (SLR) focused on the applications of ZTA. Considering the recent boom in research in this area, the conducted SLR aims to gather, and synthesize the existing studies on ZTA with a goal to provide a comprehensive taxonomy useful for researchers and practitioners who wish to apply the ZTA-oriented framework to fortify security defenses. The key contributions of this SLR are summarized as follows:
@@
  • Our study explores and synthesizes the diverse domains where ZTA has been applied, including concrete examples and case studies from the reviewed literature.
  • We present a critical analysis of the enabling technologies that facilitate ZTA adoption which can aid in understanding the technological landscape that supports ZTA frameworks.
@@ 2.1 ZTA Vs PBSM
Access control is a fundamental aspect of the field of cybersecurity, regarding the management of who or what can view, modify, or use resources in an information system. Access to information systems is prevented from unauthorized entities through two broad categories: Perimeter-Based Security Model (PBSM) and dynamic ZTA.
@@
Similarly, users can safely and remotely access the company network with Virtual Private Networks (VPN) [35]. VPNs make it more difficult for hackers to intercept or steal private information by encrypting communication between users’ devices and the company network. However, under these models, everything within the confines of a network is presumed secure and trustworthy, and that cyber threats mainly originate from external actors [14]. In such a setup, if the user has been granted access, then it could be considered trustworthy during its communication lifetime [7]. Yet, investigations indicate that insider threats are responsible for a large percentage of cyber incidents. In addition, a single account may be compromised, resulting in an escalation of privileges and presence for the attacker on the network [36] with consequences that can be severe depending on the goals of the attacker.
@@
However, the innovative approach to protect todays distributed and heterogeneous cyber landscape is Zero Trust Architecture (ZTA) [14] which is gradually gaining its foothold in the cybersecurity space. Driven by a ”never trust, always verify” philosophy, ZTA fundamentally changes how we think about network security by assuming that every part of the infrastructure is compromised and focuses on minimizing uncertainty by enforcing precise, least-privilege, per-request access decisions [5],[22]. It emphasizes situational awareness, dynamic authentication and conditional authorization, small trust zones, and continuous trust evaluation. According to a NIST special publication [37], the main goals of ZTA are data breach prevention, insider threat detection, and attacker lateral movement restriction [30]. Nowadays, the rate at which data is generated is unprecedented given the proliferation of smart devices and increasing digital transformation. This creates serious concern for both individuals and organizations when these data is disclosed or used illegally without authorization [38]. Insider threats may be malicious users, negligent users, or compromised users [21]. Because these users often have access to sensitive resources and may be well-versed in the security mechanisms in place, insider attacks pose a serious risk to organisations [39].
@@
According to a Cybersecurity Insiders report, insider attacks increased by over 47% between 2022 and 2024, affecting 34% of firms worldwide [40]. Lateral movement is a tactical exploration of a corporate network by attackers [41]. Usually, when a weak or vulnerable system (with low privileged account) is compromised, the attackers attempt to escalate their privileges while infecting other systems in the network until a set of target systems are reached [42]. The attacker continues to have access to the systems after the lateral movement in order to eventually accomplish some malicious objective, like data exfiltration or service interruption [43]. Hence, a system of continuous monitoring that examines each access request and dynamically modifies permissions based on the perceived trustworthiness of the access attempt in order to safeguard the systems, networks, and devices against malicious traffic is required [44]. This is accomplished by continuously monitoring and analysing network traffic, user and device behaviour, and other contextual data [5]. By putting security controls at the point of access, ZT actively identifies and prevents malicious or suspected individuals from gaining access to and manipulating critical data.
@@
Among the key differences between ZTA and legacy PBSM is that, while PBSM protects the enterprise (i.e., the resources) as a whole, which makes the enterprise to suffer greatly during a security breach, ZTA safeguards the individual resources [7]. It achieves this security feat by relying on some key assumptions and fundamental principles. These are contained in several articles including the study by N. Nahar et al. [14], Yuanhang et al. [18], and S. Hasan et al. [31].
@@ 2.2 ZTA Core Components
Policy Engine (PE): Within the PE, there is Zero Trust (ZT) algorithm which acts as the brain of the ZTA and makes decisions based on enterprise policies to grant, deny, or restrict access to a resource on the network [6]. The ZT algorithm also receives input from external sources such as the Continuous Diagnostic and Mitigation (CDM) systems, Records of network & system activity, activity logs, threat intelligence, etc [17], [20].
Policy Administrator (PA): The PA works closely with the PE to ultimately allow or deny a session. It sets up the PEP to enforce the decision or directive from the PE to allow or deny a session [6]. The PE together with the PA make up the control plane of the ZTA framework [18]
Policy Enforcement Point (PEP): The PEP dynamically interfaces with the entity (subject) requesting access to a resource and the resource on the network [14]. It serves as first point of contact by a request [20], and enables, monitors, and terminates access between a subject and a requested resource [29].
@@ 3 History and Publication Trend
Zero Trust Architecture (ZTA) has evolved from a security concept into an essential network security solution and is increasingly recognized by governments, industry, and academia [45]. It is a security strategy that challenges the idea that users, devices, and network components should be trusted implicitly because of where they are (i.e., their location) within a network [46]. The idea of zero trust has been around for sometime, but many articles are attributing its formal establishment to John Kindervag (2010), a research analyst working at Forrester [18], [45]. Kindervag et al. of Forrester Research [47] in a report, titled: ”No More Chewy Centers: Introducing The Zero Trust Model Of Information Security”, outlined and described four critical flaws of traditional security models that trust internal networks by default, leaving them vulnerable to insider threats and advanced attacks. In an effort to mitigate those security issues, they introduced the idea of ”Zero Trust” based on three fundamental security concepts which are: ”Ensure That All Resources Are Accessed Securely Regardless Of Location”, ”Adopt A Least Privilege Strategy And Strictly Enforce Access Control”, and ”Inspect And Log All Traffic”. The philosophy of ZTA that is currently available in the literature is based on these concepts.
@@
However, some of the earliest work on what we refer now as Zero Trust began in the early 2000s with a group of CISOs known as ”The Jericho Forum” [48], [45]. The group advocated for a new security concept called de-perimeterization, which reimagined enterprise security and arose from challenges such as managing complex firewalls amidst growing traffic volume and diversity, addressing vulnerabilities within internal network domains, and adapting to evolving business practices and working arrangements [49]. Similarly, Yong et al. [50] in 2007, while highlighting the issue of security and privacy in vehicular networks, proposed an authentication protocol using random key set. The protocol ensures user privacy based on zero-trust policy where no central entity is trusted, i.e., the vehicles do not trust public or private networks and servers. Figure  2 shows some notable events about ZTA over the years.
In the early years of ZTA, the concept received very few contributions from researchers even after the proposal of Kindervag as shown in Figure  3. The data were obtained from Scopus database using relevant keywords for research under the domain of computer science and engineering. These years can be characterized with foundational works on ZTA, defining its principles, primarily led by industry thought leaders before early research papers began exploring the architectural frameworks, practical implementations, and initial use cases in enterprise environments. It was in 2020 that the trend of ZTA-related research started to pick up where publications increased exponentially with each year almost doubling its preceding year. This surge in publications can be attributed to a confluence of factors that heightened the importance and urgency of robust cybersecurity measures which include:
@@
  • Adoption of Cloud Computing and BYOD: To effectively enable remote work, there was an increase in the adoption rate of cloud computing by organizations and businesses [52], [53] with employees increasingly using personal devices (BYOD) to access corporate systems [54]. This trend created diverse attack surfaces and underscores the need for enhanced security models that can effectively manage and secure distributed environments [55]. Zero Trust with its principles became essential for ensuring secure cloud access and mitigating risks posed by unmanaged devices and lateral movement of attackers.
  • Government Initiatives, Policy Mandates, and Standardization: Another potential that sparked publications in ZTA is governments’ initiatives and policies mandating or encouraging the adoption of Zero Trust architectures to enhance national cybersecurity resilience. For example, the executive order on cybersecurity 14028 released by the U.S. federal government in May 2021 [56]. This has led to the development of specific frameworks, such as CISA’s Zero Trust Maturity Model [57]. Similarly, the release of special publication 800 - 207 on ZTA in August 2020 by National Institute of Standards and Technology (NIST) [58] may have facilitated academic research and practical implementations, leading to increased scholarly activity.
  • Escalation of Cyber Threats and High-Profile Cyber-attacks: The period between 2020 and 2023 experienced a notable increase in cyber threats which include ransomware attacks, data breaches, etc. High-profile incidents highlight the inadequacies of traditional security approaches. These include the SolarWinds cyberattack in 2020, the Colonial pipeline ransomware attack in 2021, the Microsoft exchange server exploit in 2021, the MOVEit file transfer vulnerabilities in 2023, etc. These threats prompted organizations to adopt more resilient security architectures like Zero Trust to better protect sensitive data and critical infrastructure, and may have influence the surge in publications within this timeframe.